OpenStudy (anonymous):

write the following proportion: 8 is to 64 as 2 is to x? How do you find the answer please help

OpenStudy (anonymous):

4

OpenStudy (anonymous):

8*8=64 2*2=4 so four is the answer

OpenStudy (anonymous):

reply if I helped

OpenStudy (anonymous):

Nope \[\frac{ 8 }{ 64 }=\frac{ 2 }{ n }\] should be 16 Fraction=proportion

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so 16 sorry

OpenStudy (anonymous):

Thanks I really appreciate it

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

So if my next questions used the same numbers and says solve for x wouldnt that be the same answer as the first?

OpenStudy (anonymous):

Yes, it is the same.

OpenStudy (anonymous):

Oh wow some instuctors just try to confuse ya lol haha thanks

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

A family drove their minivan across the country, traveling from Los Angeles to Miami. The trip covered 2,950 miles and gas was $3.20 a gallon, the family paid $377.60 for gas the whole trip how many gallons of gas did they use? I got 118 by dividing $377.60/$3.20 si that right?

OpenStudy (anonymous):

it would be 118 so yes

OpenStudy (anonymous):

get a second opinion though :)

OpenStudy (anonymous):

Yes, correct!

OpenStudy (anonymous):

Great minds think alike :)

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

So how do I find 10 1/2% of 60?